# Towards Responsible Evaluation for Text-to-Speech

### v1

@@ 5.2 Challenges with Traceability and Provenance
Highly realistic TTS poses ethical challenges due to vulnerabilities to identity impersonation, deepfake abuses, and misinformation dissemination. The absence of provenance verification mechanisms erodes societal trust. It directly contradicts responsible AI principles of transparency and accountability, as illustrated by documented synthetic voice fraud Wen et al. (2025) in the financial and communication sectors. Despite these risks, current mainstream evaluations rarely assess or report verifiable markers indicating machine-generated speech, leaving TTS systems exposed to malicious misuse without adequate safeguards or traceability.
@@ 5.3 Challenges with Bias Evaluation
Bias in TTS systems emerges from multiple sources, including imbalanced training data, biased annotations, and model inductive biases. Such biases manifest as accent and gender stereotyping, as well as underrepresentation of minority speech patterns Pinhanez et al. (2024), reinforcing stereotypes and marginalizing certain communities. Despite growing attention to fairness in related AI areas, systematic bias auditing remains absent from current TTS evaluations. Prevailing evaluation practices implicitly assume universal desirability of speaker similarity or naturalness, neglecting whose voices are idealized or excluded.
@@ 5.4 Challenges with Misuse Potential and Adversarial Risk
TTS systems, especially open-source or accessible via APIs, pose growing risks of misuse and adversarial attacks Zuo et al. (2024). Synthetic speech can be exploited for impersonation, fraud, circumvention of biometric systems, or the creation of deceptive media. These risks are amplified by advances in zero-shot TTS Chen et al. (2024); Ju et al. (2024) and cross-lingual synthesis Du et al. (2024b). Despite these concerns, evaluations rarely systematically account for misuse potential. Risk assessments are often informal or absent. Existing benchmarks lack mechanisms to examine how TTS systems behave under adversarial intent, such as targeted speaker mimicry or deepfake construction.
@@ 5.5 Recommendations
To promote ethical and risk oversight in TTS evaluation, we propose the following actionable recommendations: (1) mandatory disclosure of training data provenance. Evaluation reports should move beyond ambiguous descriptors like “in-house data” by requiring detailed specifications of data sources, licensing status, and collection procedures to ensure verifiable transparency and accountability, in alignment with the EU AI Act; (2) integration of traceability indicators. We encourage the adoption of imperceptible watermarking or similar mechanisms in TTS systems. Evaluations are recommended to include metrics assessing detectability and attribution of synthetic speech; (3) construction of fairness-oriented evaluation datasets. We encourage the development of benchmarks covering diverse accents, genders, and languages to enable fair assessment of performance across underrepresented groups. (4) standardization of adversarial risk and misuse evaluation. We advocate the establishment of standardized evaluation protocols that include testing against impersonation, fraud, and other high-risk misuse scenarios.
@@ 6 Conclusion
As TTS technology continues to advance, current evaluation practices have become increasingly inadequate for capturing its full range of performance and implications. In response to this urgent need, we introduce the concept of Responsible Evaluation, structured around three progressive levels. At the first level, we advocate moving beyond conventional technical performance toward evaluation practices that faithfully and accurately reflect a model’s true capabilities, with more robust, discriminative, and comprehensive objective and subjective scoring methodologies. At the second level, we call for the adoption of standardized protocols and datasets that support meaningful comparisons and ensure high reproducibility across models and studies. At the third level, we emphasize the importance of embedding ethical and risk-aware considerations throughout the evaluation pipeline, from dataset provenance to deployment-related risks. We believe that embracing Responsible Evaluation is not only essential for advancing scientific progress in TTS but also critical for guiding TTS development in alignment with broader societal interests and responsible AI principles.
Limitations
This is a position paper that presents a conceptual and argumentative perspective aimed at advancing responsible evaluation in TTS research. Accordingly, it does not include empirical validation or implementation details. While some arguments may be controversial, they are primarily intended to provoke constructive debate and to inspire further discussion and research.
@@ Appendix A Case Study on Variants of LibriSpeech test-clean Subsets
Multiple versions of the LibriSpeech test-clean subset are used across recent TTS works, which leads to inconsistencies in reported results. One version contains 1234 utterances and is used in systems such as VALL-E Wang et al. (2023a), VALL-E 2 Chen et al. (2024), MELLE Meng et al. (2025), and PALLE Yang et al. (2025c). Another version contains 40 utterances and is used in works including NatureSpeech 3 Ju et al. (2024) and MaskGCT Wang et al. (2024b). Other subsets, such as the one used in F5-TTS Chen et al. (2025), also exist. These differences cause substantial variation in WER evaluations even for the same model.
To demonstrate this issue, we evaluate the open-sourced MaskGCT [^1] https://huggingface.co/amphion/MaskGCT on two commonly used variants of the test-clean subset. WER is computed between ASR transcription of synthesized audio and the ground-truth text, using the HuBERT-Large ASR model [^2] https://huggingface.co/facebook/hubert-large-ls960-ft Hsu et al. (2021). The WER differs significantly across the two versions, ranging from 2.63 to 4.22, as shown in Table 1. This observation argues the importance of clearly reporting dataset versions and evaluation protocols to ensure fair and reproducible comparisons.
@@
40 utterances Wang et al. (2024b) 2.63
1234 utterances Yang et al. (2025c) 4.22
@@ Appendix B Case Study on Inconsistencies in SIM-o Evaluation Protocols
SIM-o is defined as the cosine similarity between speaker embeddings extracted from original speech and synthesized speech. Commonly, SIM-o is computed using WavLM-TDNN [^3] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification#pre-trained-models Chen et al. (2022), where the score ranges within [1,1][-1,1], with higher values indicating greater speaker similarity.
However, there are two practices for computing SIM-o for the continuation task. One approach, adopted by VALL-E Wang et al. (2023a), computes speaker similarity between the first 3-second ground-truth speech prompt and the remaining synthesized speech, excluding the prompt. Alternatively, another approach, as used in VALL-E 2 Chen et al. (2024), computes the similarity between the full synthesized speech, including the prompt and the entire ground-truth speech.
@@
Without Prompt Wang et al. (2023a) 0.754
With Prompt Chen et al. (2024) 0.905
@@ Appendix C Comparison between WER and Perceived Intelligibility
To examine the relationship between WER and perceptual intelligibility, we conduct an analysis. We perform ASR on both synthesized speech from MELLE Meng et al. (2025) and ground-truth speech using the Conformer-Transducer model [^4] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge. We recruit 10 graduate students with research experience in TTS as raters and conduct a MOS test focusing on perceived intelligibility, denoted as WER-MOS. Subjective ratings are computed by manually transcribing each sample and calculating its corresponding WER.
@@
MELLE Meng et al. (2025) 1.47 1.05
Appendix D Alternative Views
Our perspectives are intended to initiate ongoing discussion. While acknowledging diverse views and potential curiosities, we objectively examine several alternate viewpoints:
Concerns about increased evaluation complexity
Some researchers and practitioners caution that introducing additional evaluation metrics could complicate the evaluation process, particularly in industrial contexts where scalability and efficiency are critical. They also note that an overabundance of criteria might risk fragmenting TTS evaluation practices, thereby reducing comparability and standardization.
Response While expanding evaluation dimensions and introducing new metrics may pose short-term challenges, such efforts are essential to ensure that TTS evaluation evolves in step with technological advances and real-world requirements. As in many areas of technology, development often moves from diversification to convergence, ultimately leading to more unified and stable practices.
Balancing rapid progress with legal and ethical considerations
Some researchers and practitioners caution that excessive emphasis on legal and ethical aspects could inadvertently slow technological innovation. Especially, overly restrictive interpretations of data copyright may constrain progress in low-resource languages and domains where available data are scarce.
Response We acknowledge that in low-resource settings, limited copyright awareness and the scarcity of high-quality data genuinely present challenges to TTS development. However, these challenges are not insurmountable. Doctrines such as Fair Use provide avenues for ethically grounded data utilization, and techniques such as few-shot learning can reduce reliance on large-scale datasets. Together, these approaches offer a responsible path toward sustainable TTS advancement.

### v2

@@ 5.2 Challenges with Traceability and Provenance
Highly realistic TTS poses ethical challenges due to vulnerabilities to misinformation dissemination (Liu et al., 2024). The absence of provenance verification mechanisms erodes societal trust. It directly contradicts responsible AI principles of transparency and accountability, as illustrated by documented synthetic voice fraud (Wen et al., 2025) in the financial and communication sectors. Despite these risks, current mainstream evaluations rarely report verifiable markers indicating machine-generated speech, leaving TTS systems exposed to malicious misuse without adequate safeguards or traceability.
@@ 5.3 Challenges with Bias Evaluation
Bias in TTS systems emerges from multiple sources, including imbalanced training data, biased annotations, and model inductive biases. Such biases manifest as accent and gender stereotyping, as well as underrepresentation of minority speech patterns (Pinhanez et al., 2024), reinforcing stereotypes and marginalizing certain communities. Despite growing attention to fairness in related AI areas, systematic bias auditing remains absent from current TTS evaluations. Prevailing evaluation practices implicitly assume universal desirability of speaker similarity or naturalness, neglecting whose voices are idealized or excluded.
@@ 5.4 Challenges with Misuse Potential and Adversarial Risk
The increasing accessibility of TTS systems, particularly open-source models and API-based services, poses growing risks of misuse and adversarial attacks (Zuo et al., 2024). Synthetic speech can be exploited for impersonation, fraud, circumvention of biometric authentication systems (Wang et al., 2024b), or the creation of deceptive media. In addition, adversarial TTS attacks can generate audio that sounds natural to humans yet triggers unintended actions by voice-controlled assistants such as Alexa or Siri, enabling malicious commands. Despite these concerns, evaluations rarely systematically account for misuse potential. Risk assessments are often informal or absent. Existing evaluations lack mechanisms to examine how TTS systems behave under adversarial intent, such as targeted speaker mimicry or deepfake construction.
@@ 5.5 Recommendations
To promote ethical and risk oversight in TTS evaluation, we propose the following actionable recommendations: (1) mandatory disclosure of training data provenance. Evaluation reports should move beyond ambiguous descriptors like “in-house data” by requiring detailed specifications of data sources, licensing status, and collection procedures to ensure verifiable transparency and accountability, in alignment with the EU AI Act. (2) integration of traceability indicators. We encourage the adoption of imperceptible watermarking or similar mechanisms in TTS systems. Evaluations are recommended to include metrics assessing detectability and attribution of synthetic speech. (3) construction of fairness-oriented evaluation datasets. We encourage the development of benchmarks covering diverse accents, genders, and languages to enable fair assessment of performance across underrepresented groups. (4) standardization of adversarial risk and misuse evaluation. We advocate establishing standardized evaluation protocols that include testing against impersonation, fraud, and other high-risk misuse scenarios.
6 Alternative Views
Our perspectives are intended to stimulate further discussion. While we acknowledge diverse viewpoints, we discuss several alternative views to our position below:
Alternative View 1: Concerns about increased evaluation complexity
Some practitioners caution that introducing additional evaluation metrics could complicate the evaluation process, particularly in industrial contexts where scalability and efficiency are critical. They also note that an overabundance of criteria might risk fragmenting TTS evaluation practices, thereby reducing comparability and standardization. We believe that while expanding evaluation dimensions and introducing new metrics may pose short-term challenges, such efforts are essential to ensure that TTS evaluation evolves in step with technological advances and real-world requirements. As in many areas of technology, development often moves from diversification to convergence, ultimately leading to unified and stable practices.
Alternative View 2: Balancing rapid progress with legal and ethical considerations
Some practitioners caution that excessive emphasis on legal and ethical aspects could inadvertently slow technological innovation. Especially, overly restrictive interpretations of data copyright may constrain progress in low-resource languages and domains where available data are scarce. We acknowledge that in low-resource settings, limited copyright awareness and the scarcity of high-quality data genuinely present challenges to TTS development. However, these challenges are not insurmountable. Doctrines such as Fair Use provide avenues for ethically grounded data utilization, and techniques such as few-shot learning can reduce reliance on large-scale datasets. Together, these approaches offer a responsible path toward sustainable TTS advancement.
@@ 7 Conclusion
As TTS technology continues to advance, current evaluation practices have become increasingly inadequate for capturing its full range of capabilities and implications. In response to this urgent need, we introduce the concept of Responsible Evaluation, structured around three progressive levels. At the first level, we advocate moving beyond conventional technical performance toward evaluation practices that faithfully and accurately reflect a model’s true capabilities, with more robust, discriminative, and comprehensive objective and subjective scoring methodologies. At the second level, we call for the adoption of standardized protocols and datasets that support fair comparisons and ensure reproducibility across models and studies. At the third level, we emphasize the importance of integrating ethical and risk-aware considerations throughout the evaluation pipeline, from dataset provenance to deployment-related risks. We believe that embracing Responsible Evaluation is not only essential for advancing scientific progress in TTS but also critical for guiding TTS development in alignment with broader societal interests and responsible AI principles.
Impact Statement
This is a position paper that presents a conceptual and argumentative perspective aimed at advancing Responsible Evaluation in TTS research. While some arguments may be controversial, they are primarily intended to provoke constructive debate and to inspire further discussion and research. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Case Study on Variants of LibriSpeech test-clean Subsets
Multiple versions of the LibriSpeech test-clean subset are used across recent TTS works, which leads to inconsistencies in reported results. One version contains 1234 utterances and is used in systems such as VALL-E (Wang et al., 2023), VALL-E 2 (Chen et al., 2024), MELLE (Meng et al., 2025), and PALLE (Yang et al., 2025b). Another version contains 40 utterances and is used in works including NatureSpeech 3 (Ju et al., 2024) and MaskGCT (Wang et al., 2024c). Other subsets, such as the one used in F5-TTS (Chen et al., 2025), also exist. These differences cause substantial variation in WER evaluations even for the same model.
To demonstrate this issue, we evaluate the open-sourced MaskGCT [^1] https://huggingface.co/amphion/MaskGCT on two commonly used variants of the test-clean subset. WER is computed between ASR transcription of synthesized audio and the ground-truth text, using the HuBERT-Large ASR model [^2] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021). The WER differs significantly across the two versions, ranging from 2.63 to 4.22, as shown in Table 1. This observation argues the importance of clearly reporting dataset versions and evaluation protocols to ensure fair and reproducible comparisons.
@@
40 utterances (Wang et al., 2024c) 2.63
1234 utterances (Yang et al., 2025b) 4.22
@@ Appendix B Case Study on Inconsistencies in SIM-o Evaluation Protocols
SIM-o is defined as the cosine similarity between speaker embeddings extracted from original speech and synthesized speech. Commonly, SIM-o is computed using WavLM-TDNN [^3] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification#pre-trained-models (Chen et al., 2022), where the score ranges within [1,1][-1,1], with higher values indicating greater speaker similarity.
However, there are two practices for computing SIM-o for the continuation task. One approach, adopted by VALL-E (Wang et al., 2023), computes speaker similarity between the first 3-second ground-truth speech prompt and the remaining synthesized speech, excluding the prompt. Alternatively, another approach, as used in VALL-E 2 (Chen et al., 2024), computes the similarity between the full synthesized speech, including the prompt and the entire ground-truth speech.
@@
Without Prompt (Wang et al., 2023) 0.754
With Prompt (Chen et al., 2024) 0.905
@@ Appendix C Comparison between WER and Perceived Intelligibility
To examine the relationship between WER and perceptual intelligibility, we conduct an analysis. We perform ASR on both synthesized speech from MELLE (Meng et al., 2025) and ground-truth speech using the Conformer-Transducer model [^4] https://huggingface.co/nvidia/stt_en_conformer_transducer_xlarge. We recruit 10 graduate students with research experience in TTS as raters and conduct a MOS test focusing on perceived intelligibility, denoted as WER-MOS. Subjective ratings are computed by manually transcribing each sample and calculating its corresponding WER.
@@
MELLE (Meng et al., 2025) 1.47 1.06